Convert 0.8 to the common fraction in its simple form. Show all necessary working.


1
Expert's answer
2021-04-14T00:58:07-0400

0.8 that means 8 / 10. It is a simple form of a number but not finally. We will shorten simple form to get finally form. 8 = 4 * 2 and 10 = 5 * 2 => (4 * 2) / (5 * 2) => After this we will short number 2.

So finally we will get 4 / 5 it is finally simple from.